OpenStudy (anonymous):

The temperature of a chemical reaction ranges between 20°C and 160°C. The temperature is at its lowest point when t = 0, the temperature is 90 degrees and increasing. What is a sine function that would model this reaction?

OpenStudy (anonymous):

possible answers: f(x) = 90 sin π/4 x + 70 f(x) = 70 sin π/4 x + 90 f(x) = 70 sin 8x + 90 f(x) = 90 sin 8x + 70